OpenStudy (anonymous):

find real numbers a and b such that a) a+bi = 4 b) (1-2i) (a+bi) = -5-10i c) (a+2i) (1+bi) = 17-19i

OpenStudy (mr.math):

\(a+bi=c+di \iff a=c \text{ and } b=d.\)

OpenStudy (mr.math):

a) \(a=4, b=0\)

OpenStudy (mr.math):

b)\((1-2i)(a+bi)=a+bi-2ai+2b=(a+2b)+i(b-2a)=-5-10i\) \(\iff a+2b=-5 \text{ and }b-2a=-10 \implies a=3, b=-4\)

OpenStudy (mr.math):

You try the last problem.

OpenStudy (anonymous):

i will try!!! :P can you just check if am doing it right then?

OpenStudy (mr.math):

Of course!